# The Racah algebra of rank 2: Properties, symmetries and representation

### v1

@@ 4.1 The rank-1 Racah algebra
Let us recall some of the representation theory of the rank-one universal Racah algebra. This representation theory has been studied by several authors [3, 13] but here we focus on the theory explained by Huang and Bockting-Conrad [13]. As discussed in the previous section, we make a choice of basis using linealry independent generators $A=C_{23}$ and $B=C_{12}.$ The relations are then:
$~{}[A,B]$ $=$ $2D,$ (4.1)
$~{}[A,D]$ $=$ $\{A,B\}+A^{2}-\delta A+\alpha,$ (4.2)
$~{}[D,B]$ $=$ $\{A,B\}+B^{2}-\delta B-\beta,$ (4.3)
@@
$A=C_{23},$ $\alpha=(C_{2}-C_{3})(C_{1}-C_{123}),$ (4.4)
$B=C_{12},$ $\beta=(C_{1}-C_{2})(C_{3}-C_{123}),$ (4.5)
$D=D_{123},$ $\delta=C_{123}+C_{1}+C_{2}+C_{3}.$ (4.6)
The authors give a infinite dimensional $R(3)$ module (Prop. 3.1, [13]) and later show that this module is isomorphic to the quotient of the universal enveloping algebra by an ideal, similar to the construction of Verma modules for Lie algebra representations.
In this case, the action of $A$ and $B$ on states $|j\rangle$ in the $R(3)$-module are as follow :
@@
where the raising coefficient of $A$ has been normalized to 1. In other words, $A$ acts centrally and east, while $B$ acts centrally and west, considering that the states are ordered such that $|j+1\rangle$ is east of $|j\rangle$.
@@
$C_{1}=c_{1}(c_{1}-1),\qquad C_{2}=c_{2}(c_{2}-1),\qquad C_{3}=c_{3}(c_{3}-1),$ (4.9)
$C_{123}=(c_{1}+c_{2}+c_{3}+N)(c_{1}+c_{2}+c_{3}+N+1).$ (4.10)
Note that in the extension to rank-two, $C_{123}$ will no longer be central. However, this Racah algebra of rank 1 will appear as a Racah subalgebra in the rank-2 case.
@@ 4.2 The rank-2 Racah algebra
In the remainder of the paper, we make a choice of basis for our module, the contiguous basis, breaking both the permutation symmetry of the original tensor product representation and the dihedral symmetry of the previous section. However, it is possible to use the $\mathcal{D}_{5}$ and $\mathcal{P}_{4}$ transformation to create a representation in terms of other generators.
@@
and we express the central terms using constants $c_{i}$, i.e.
$\omega_{0}=C_{1234}=c_{0}(c_{0}-1),\qquad\omega_{i}=C_{i}=c_{i}(c_{i}-1),\qquad i=1,...,4.$ (4.11)
@@
$\Gamma_{0}=\dfrac{1}{2}[C_{123},C_{234}],\qquad\Gamma_{1}=\dfrac{1}{2}[C_{234},C_{12}],\qquad\Gamma_{2}=\dfrac{1}{2}[C_{12},C_{23}],$
$\Gamma_{3}=\dfrac{1}{2}[C_{23},C_{34}],\qquad\Gamma_{4}=\dfrac{1}{2}[C_{34},C_{123}].$ (4.12)
Notice that $\Omega_{0}$ and $\Omega_{4}$ form an $R(3)$ subalgebra together with $\Omega_{2}$, so we extend the notation from section 4.1, the rank-one case, to construct the representation. We assume that $\Omega_{0}$ and $\Omega_{4}$ act similarly to $A$ and $B$ in (4.7,4.8), that is
$C_{12}|t,s\rangle=\varphi_{t,s}|t-1,s\rangle+\theta_{t,s}|t,s\rangle,\qquad C_{23}|t,s\rangle=\theta^{*}_{t,s}|t,s\rangle+|t+1,s\rangle,$ (4.13)
where the coefficient of the raising term of $C_{23}$ is normalized to 1. We do not make any additional assumptions on $\Omega_{1}$ $\Omega_{2}$ and $\Omega_{3}$ except if applied on a state, they send a state to a linear combination of itself and its 8 closest neighbours, e.g.
$C_{123}|t,s\rangle$ $=$ $\hat{\mu}_{t,s}|t-1,s+1\rangle+\hat{\nu}_{t,s}|t,s+1\rangle+\hat{\xi}_{t,s}|t+1,s+1\rangle$ (4.14)
$+\mu_{t,s}|t-1,s\rangle+\nu_{t,s}|t,s\rangle+\xi_{t,s}|t+1,s\rangle$
$+\check{\mu}_{t,s}|t-1,s-1\rangle+\check{\nu}_{t,s}|t,s-1\rangle+\check{\xi}_{t,s}|t+1,s-1\rangle.$
Since $C_{12}$, $C_{23}$ and $C_{123}$ are the non-central generator of a rank-1 Racah subalgebra, one can use the presentation equations of the rank-1 Racah (4.14.3) to determine the coefficients. Considering that $C_{12}$ and $C_{23}$ can only move a state east-west, this implies that $C_{123}$ can only move east-west. Furthermore, by solving all the coefficients so they satisfy (4.14.3), one gets that $C_{123}$ is forced to be central, that is
$C_{123}|t,s\rangle$ $=$ $\nu_{t,s}|t,s\rangle.$ (4.15)
@@
$C_{34}|t,s\rangle$ $=$ $\hat{\phi}^{*}_{t,s}|t-1,s+1\rangle+\hat{\vartheta}^{*}_{t,s}|t,s+1\rangle+\hat{\psi}^{*}_{t,s}|t+1,s+1\rangle$ (4.16)
$+\phi^{*}_{t,s}|t-1,s\rangle+\vartheta^{*}_{t,s}|t,s\rangle+\psi^{*}_{t,s}|t+1,s\rangle$
$+\check{\phi}^{*}_{t,s}|t-1,s-1\rangle+\check{\vartheta}^{*}_{t,s}|t,s-1\rangle+\check{\psi}^{*}_{t,s}|t+1,s-1\rangle,$
$C_{234}|t,s\rangle$ $=$ $\hat{\phi}_{t,s}|t-1,s+1\rangle+\hat{\vartheta}_{t,s}|t,s+1\rangle+\hat{\psi}_{t,s}|t+1,s+1\rangle$ (4.17)
$+\phi_{t,s}|t-1,s\rangle+\vartheta_{t,s}|t,s\rangle+\psi_{t,s}|t+1,s\rangle$
$+\check{\phi}_{t,s}|t-1,s-1\rangle+\check{\vartheta}_{t,s}|t,s-1\rangle+\check{\psi}_{t,s}|t+1,s-1\rangle,$
@@
$C_{34}|t,s\rangle$ $=$ $\hat{\phi}^{*}_{t,s}|t-1,s+1\rangle+\hat{\vartheta}^{*}_{t,s}|t,s+1\rangle$ (4.18)
$+\phi^{*}_{t,s}|t-1,s\rangle+\vartheta^{*}_{t,s}|t,s\rangle$
@@
$C_{234}|t,s\rangle$ $=$ $\hat{\vartheta}_{t,s}|t,s+1\rangle+\hat{\psi}_{t,s}|t+1,s+1\rangle$ (4.19)
$+\vartheta_{t,s}|t,s\rangle+\psi_{t,s}|t+1,s\rangle$
@@
Solving all the recurrence equations is quite a task, so we will not show all the calculations. However, it is convenient to first solve the Racah subalgebra involving $C_{12},C_{23}$ and $C_{123}$, i.e. using the commuting and the two inner relations, and then continue with the other ones. The coefficients can be found in the appendix B. To illustrate the direction of the action and the meaning of each coefficient, we produced the following diagram for each generators. The other elements of the rank-2 Racah algebra can be constructed using equation (2.12).
@@
Figure 4: Direction of the action of each non-diagonal generator with the dependency on the parameter $t$ and $s$.
@@ 5 Conclusion
First, we have given a “universal” definition of the Racah algebra in terms of generators and relations that can be seen as an extension of the rank-1 Racah algebra using sets instead of indices. The main properties and relations generating $R(4)$ and $R(n)$ are summarized in Theorem 1 and Theorem 2 respectively. Additional properties can be found throughout Section 2. Most significantly, we obtain the commutation relations for general $P_{ij}$ and $D_{jk\ell}$ as well as higher-order commutation relations, for example (2.25), given in other definitions[5]. Similarly, we obtain the identity (2.29) first identified in [5] as arising from associativity in the algebra. We further show that in $R(4)$ no additional relations arise from the Jacobi identities in the algebra. This result is important for our investigation into the representation theory of $R(4)$ and in particular for future work describing Verma modules and PBW-type bases for the algebra, following on similar work as in [2, 13, 14] and conjuectures in [7].
[…]
@@
Finally, we provide a novel representation for $R(4)$ using a split basis, an extension of the representation by Huang and Bockting-Conrad [13]. The explicit coefficients from the action of the generators can be found in Appendix B. A representation of $R(4)$ already exists [4] by Crampé, Frappat and Ragoucy. However that representation and our representation differ. The main difference resides in the fact that the representation by Crampé et al. requires that 2 generators are diagonalizable, while ours only requires one generator to be diagonalizable, which is a consequence of our choice of a split basis. In addition, Crampé et al. assumed that their representation are from the so-called special Racah, while we did not impose this constraint.
Acknowlegment
The authors (SB) would like to thank Nicolas Crampé and Éric Ragoucy for the fruitful discussion on the subject.
SP would like to acknowledge the Simons Foundation Collaboration grant #3192112 as well as support as a CRM-Simons professorship that she held while this research was being conducted.
Appendix A Appendix — Proofs of double commutators
@@
Before we begin the computations, a bit of discussion of the strategy. In each case, we try to reduce the equations to triples of P’s and in the case that a term contains commuting P’s we try to move that pair to the left. In line with this, we move inner relations to the right and outer relations to the left. It will also be useful to the following forms of the inner relation (2.32).
$[P_{ij},D_{ijk}]=P_{ij}(P_{jk}-P_{ik})+2D_{ijk}-2P_{i}P_{jk}+2P_{j}P_{ik}.$
@@
A.1 Case $\{P_{i\ell},D_{ijk},D_{jk\ell}\}.$
We begin with the Jacobi relation for the set $\{P_{i\ell},D_{ijk},D_{jk\ell}\}.$
$[P_{i\ell},[D_{ijk},D_{jkl}]]$ $=$ $[P_{i\ell},P_{jk}(D_{ji\ell}+D_{ki\ell})]$
$=$ $P_{jk}\left(P_{i\ell}(P_{j\ell}+P_{k\ell}-P_{ij}-P_{ik})+2D_{i\ell j}+2D_{i\ell k}\right)$
$-2P_{i}P_{jk}(P_{\ell j}+P_{\ell k})+2P_{\ell}P_{jk}(P_{ij}+P_{ik})$
$=$ $P_{jk}P_{i\ell}(P_{j\ell}+P_{k\ell}-P_{ij}-P_{ik})$
$-2P_{jk}(D_{ji\ell}+D_{ki\ell})-2P_{i}P_{jk}\left(P_{j\ell}+P_{k\ell}\right)+2P_{\ell}P_{jk}\left(P_{ij}+P_{jk}\right).$
@@
$[D_{ijk},[D_{jk\ell},P_{i\ell}]]$ $=$ $[[P_{i\ell},D_{\ell jk}],D_{ijk}]$
$=$ $[P_{ik}P_{\ell j}-P_{\ell k}P_{ij},D_{ijk}]$
$=$ $[P_{\ell j}P_{ik}-P_{\ell k}P_{ij},D_{ijk}]$
$=$ $[P_{\ell j},D_{jki}]P_{ik}-[P_{\ell k},D_{kij}]P_{ij}+P_{\ell j}[P_{ki},D_{kij}]-P_{\ell k}[P_{ij},D_{ijk}]$
$=$ $(P_{i\ell}P_{jk}-P_{ij}P_{kl})P_{ik}-(P_{\ell j}P_{ik}-P_{jk}P_{i\ell})P_{ij}$
$+P_{\ell j}\left(P_{ki}(P_{ij}-P_{jk})-2D_{ijk}-2P_{k}P_{ij}+2P_{i}P_{jk}\right)$
$-P_{\ell k}\left(P_{ij}(P_{jk}-P_{ki})-2D_{ijk}-2P_{i}P_{jk}+2P_{j}P_{ki}\right)$
$=$ $P_{i\ell}P_{jk}(P_{ik}+P_{ij})-(P_{ik}P_{j\ell}+P_{ij}P_{k\ell})P_{jk}$
$+2(P_{k\ell}-P_{j\ell})D_{ijk}+2P_{i}(P_{j\ell}+P_{k\ell})P_{jk}-2P_{j}P_{k\ell}P_{ik}-2P_{k}P_{\ell j}P_{ij}.$

### v2

@@ 4.1 The rank-1 Racah algebra
Let us recall some of the representation theory of the rank-one universal Racah algebra. This representation theory has been studied by several authors but here we focus on the theory explained by Huang and Bockting-Conrad [11]. As discussed in the previous section, we make a choice of basis using linearly independent generators $A=C_{23}$ and $B=C_{12}$. The relations are then
$[A,B]=2D,\qquad[A,D]=\{A,B\}+A^{2}-\delta A+\alpha,$ (4.1)
$[D,B]=\{A,B\}+B^{2}-\delta B-\beta,$ (4.2)
@@
$A=C_{23},\qquad$ $\alpha=(C_{2}-C_{3})(C_{1}-C_{123}),\qquad$ $B=C_{12},\qquad$ $\beta=(C_{1}-C_{2})(C_{3}-C_{123}),$
$D=D_{123},\qquad$ $\delta=C_{123}+C_{1}+C_{2}+C_{3}.$
The authors give an infinite-dimensional $R(3)$ module [11, Proposition 3.1] and later show that this module is isomorphic to the quotient of the universal enveloping algebra by an ideal, similar to the construction of Verma modules for Lie algebra representations.
In this case, the action of $A$ and $B$ on states $|j\rangle$ in the $R(3)$-module are as follows:
@@
where the raising coefficient of $A$ has been normalized to 1. In other words, $A$ acts centrally and east, while $B$ acts centrally and west, considering that the states are ordered such that $|j+1\rangle$ is east of $|j\rangle$. The coefficients $\theta_{j}$, $\theta_{j}^{*}$ and $\varphi_{j}$ can be found in the paper [11] or by setting the extra parameter $s$ to zero in Appendix B.
@@
$C_{1}=c_{1}(c_{1}-1),\qquad C_{2}=c_{2}(c_{2}-1),\qquad C_{3}=c_{3}(c_{3}-1),$
$C_{123}=(c_{1}+c_{2}+c_{3}+N)(c_{1}+c_{2}+c_{3}+N+1).$
Note that in the extension to rank-2, $C_{123}$ will no longer be central. However, this Racah algebra of rank 1 will appear as a Racah subalgebra in the rank-2 case.
@@ 4.2 The rank-2 Racah algebra
In the remainder of the paper, we make a choice of basis for our module, the contiguous basis, breaking both the permutation symmetry of the original tensor product representation and the dihedral symmetry of the previous section. However, it is possible to use the $\mathcal{D}_{5}$ and $\mathcal{P}_{4}$ transformations to create a representation in terms of other generators.
@@
and we express the central terms using constants $c_{i}$, i.e.,
$\omega_{0}=C_{1234}=c_{0}(c_{0}-1),\qquad\omega_{i}=C_{i}=c_{i}(c_{i}-1),\qquad i=1,\dots,4.$
@@
$\Gamma_{0}=\tfrac{1}{2}[C_{123},C_{234}],\qquad$ $\Gamma_{1}=\tfrac{1}{2}[C_{234},C_{12}],\qquad$ $\Gamma_{2}=\tfrac{1}{2}[C_{12},C_{23}],$
$\Gamma_{3}=\tfrac{1}{2}[C_{23},C_{34}],\qquad$ $\Gamma_{4}=\tfrac{1}{2}[C_{34},C_{123}].$
Notice that $\Omega_{0}$ and $\Omega_{4}$ form an $R(3)$ subalgebra together with $\Omega_{2}$, so we extend the notation from Section 4.1, the rank-1 case, to construct the representation. We assume that $\Omega_{0}$ and $\Omega_{4}$ act similarly to $A$ and $B$ in (4.3), (4.4), that is $C_{12}|t,s\rangle=\varphi_{t,s}|t-1,s\rangle+\theta_{t,s}|t,s\rangle$, $C_{23}|t,s\rangle=\theta^{*}_{t,s}|t,s\rangle+|t+1,s\rangle$, where the coefficient of the raising term of $C_{23}$ is normalized to 1. We do not make any additional assumptions on $\Omega_{1}$, $\Omega_{2}$ and $\Omega_{3}$ except if applied on a state, they send a state to a linear combination of itself and its 8 closest neighbours, e.g.,
$C_{123}|t,s\rangle=\hat{\mu}_{t,s}|t-1,s+1\rangle+\hat{\nu}_{t,s}|t,s+1\rangle+\hat{\xi}_{t,s}|t+1,s+1\rangle+\mu_{t,s}|t-1,s\rangle$
${}+\nu_{t,s}|t,s\rangle+\xi_{t,s}|t+1,s\rangle+\check{\mu}_{t,s}|t-1,s-1\rangle+\check{\nu}_{t,s}|t,s-1\rangle+\check{\xi}_{t,s}|t+1,s-1\rangle.$
Since $C_{12}$, $C_{23}$ and $C_{123}$ are the non-central generator of a rank-1 Racah subalgebra, one can use the presentation equations of the rank-1 Racah (4.1)–(4.2) to determine the coefficients. Considering that $C_{12}$ and $C_{23}$ can only move a state east-west, this implies that $C_{123}$ can only move east-west. Furthermore, by solving all the coefficients so they satisfy (4.1)–(4.2), one gets that $C_{123}$ is forced to be central, that is $C_{123}|t,s\rangle=\nu_{t,s}|t,s\rangle$.
@@
$C_{34}|t,s\rangle=\hat{\phi}^{*}_{t,s}|t-1,s+1\rangle+\hat{\vartheta}^{*}_{t,s}|t,s+1\rangle+\hat{\psi}^{*}_{t,s}|t+1,s+1\rangle+\phi^{*}_{t,s}|t-1,s\rangle+\vartheta^{*}_{t,s}|t,s\rangle$
${}+\psi^{*}_{t,s}|t+1,s\rangle+\check{\phi}^{*}_{t,s}|t-1,s-1\rangle+\check{\vartheta}^{*}_{t,s}|t,s-1\rangle+\check{\psi}^{*}_{t,s}|t+1,s-1\rangle,$
$C_{234}|t,s\rangle=\hat{\phi}_{t,s}|t-1,s+1\rangle+\hat{\vartheta}_{t,s}|t,s+1\rangle+\hat{\psi}_{t,s}|t+1,s+1\rangle+\phi_{t,s}|t-1,s\rangle+\vartheta_{t,s}|t,s\rangle$
${}+\psi_{t,s}|t+1,s\rangle+\check{\phi}_{t,s}|t-1,s-1\rangle+\check{\vartheta}_{t,s}|t,s-1\rangle+\check{\psi}_{t,s}|t+1,s-1\rangle.$
@@
$C_{34}|t,s\rangle=\hat{\phi}^{*}_{t,s}|t-1,s+1\rangle+\hat{\vartheta}^{*}_{t,s}|t,s+1\rangle+\phi^{*}_{t,s}|t-1,s\rangle+\vartheta^{*}_{t,s}|t,s\rangle$
@@
$C_{234}|t,s\rangle=\hat{\vartheta}_{t,s}|t,s+1\rangle+\hat{\psi}_{t,s}|t+1,s+1\rangle+\vartheta_{t,s}|t,s\rangle+\psi_{t,s}|t+1,s\rangle$
${}+\check{\vartheta}_{t,s}|t,s-1\rangle+\check{\psi}_{t,s}|t+1,s-1\rangle.$
@@
Solving all the recurrence equations is quite a task, so we will not show all the calculations. However, it is convenient to first solve the Racah subalgebra involving $C_{12},C_{23}$ and $C_{123}$, i.e., using the commuting and the two inner relations, and then continue with the other ones. The coefficients can be found in Appendix B. To illustrate the direction of the action and the meaning of each coefficient, we produced the following diagram for each generators. The other elements of the rank-2 Racah algebra can be constructed using equation (2.9).
@@
Figure 4: Direction of the action of each non-diagonal generator with the dependency on the parameters $t$ and $s$.
@@ 5 Conclusion
First, we have given a “universal” definition of the Racah algebra in terms of generators and relations that can be seen as an extension of the rank-1 Racah algebra using sets instead of indices. The main properties and relations generating $R(4)$ and $R(n)$ are summarized in Theorems 2.11 and 2.12, respectively. Additional properties can be found throughout Section 2. Most significantly, we obtain the commutation relations for general $P_{ij}$ and $D_{jk\ell}$ as well as higher-order commutation relations, for example, (2.16), given in other definitions [4]. Similarly, we obtain the identity (2.18) first identified in [4] as arising from associativity in the algebra. We further show that in $R(4)$ no additional relations arise from the Jacobi identities in the algebra. This result is important for our investigation into the representation theory of $R(4)$ and in particular for future work describing Verma modules and PBW-type bases for the algebra, following on similar work as in [2, 11, 12] and conjectures in [5].
Next, we have explored the symmetries of some $R(4)$ generators, that is the contiguous basis. While this basis is unique, that is the set of generators is composed of operators with continuous indices, i.e., $\{C_{1},C_{2},C_{3},C_{4},C_{12},C_{23},C_{34},C_{123},C_{234},C_{1234}\}$, many other sets of generators are isomorphic. By using the permutation symmetries $\mathcal{P}_{4}$ on the indices, one can get another set of generators with the same properties and commutation relations. With this symmetry, one can make a partial abstraction of the choice of generators to create a presentation and a representation. That special set of basis possesses an addition symmetry group, which is an automorphism of the set, that is the dihedral group $\mathcal{D}_{5}$. Indeed, the symmetry of the generators can be illustrated using a pentagon, as shown in Figure 3. This choice of basis leads to elegant properties. Among others, the rank-1 Racah algebra $R(3)$ using the indices $\{1,2,3\}$ is a subalgebra of $R(4)$, and using the $\mathcal{D}_{5}$ symmetry, 4 additional rank-1 Racah subalgebras arise naturally. Furthermore, if one apply $\mathcal{D}_{5}$ symmetries to the Casimir operator of $R(3)$, then 4 other Casimir operators arise.
@@
Finally, we provide a novel representation for $R(4)$ using a split basis, an extension of the representation by Huang and Bockting-Conrad [11]. The explicit coefficients from the action of the generators can be found in Appendix B. A representation of $R(4)$ already exists [3] by Crampé, Frappat and Ragoucy. However that representation and our representation differ. The main difference resides in the fact that the representation by Crampé et al. requires that 2 generators are diagonalizable, while ours only requires one generator to be diagonalizable, which is a consequence of our choice of a split basis. In addition, Crampé et al. assumed that their representation is from the so-called special Racah, while we did not impose this constraint. However, using our representation, we calculated the values of the Casimir operators $\mathfrak{C}_{i}$${i=0,\dots,4}$, and we obtained that all of them vanish. Either our hypothesis led us to assume implicitly special Racah, or there are no other cases, that is the so-called special Racah is in fact Racah itself.
Appendix A Proofs of double commutators
@@
Before we begin the computations, a bit of discussion of the strategy. In each case, we try to reduce the equations to triples of $P$’s and in the case that a term contains commuting $P$’s we try to move that pair to the left. In line with this, we move inner relations to the right and outer relations to the left. It will also be useful to the following forms of the inner relation (2.20):
$[P_{ij},D_{ijk}]=P_{ij}(P_{jk}-P_{ik})+2D_{ijk}-2P_{i}P_{jk}+2P_{j}P_{ik},$
@@
A.1 Case $\boldsymbol{\{P_{i\ell},D_{ijk},D_{jk\ell}\}}$
We begin with the Jacobi relation for the set $\{P_{i\ell},D_{ijk},D_{jk\ell}\}$
$[P_{i\ell},[D_{ijk},D_{jkl}]]$
$\qquad=[P_{i\ell},P_{jk}(D_{ji\ell}+D_{ki\ell})]$
$\qquad=P_{jk}(P_{i\ell}(P_{j\ell}+P_{k\ell}-P_{ij}-P_{ik})+2D_{i\ell j}+2D_{i\ell k})-2P_{i}P_{jk}(P_{\ell j}+P_{\ell k})$
$\phantom{\qquad=}{}+2P_{\ell}P_{jk}(P_{ij}+P_{ik})$
$\qquad=P_{jk}P_{i\ell}(P_{j\ell}+P_{k\ell}-P_{ij}-P_{ik})-2P_{jk}(D_{ji\ell}+D_{ki\ell})-2P_{i}P_{jk}(P_{j\ell}+P_{k\ell})$
$\phantom{\qquad=}{}+2P_{\ell}P_{jk}(P_{ij}+P_{jk}).$ (A.1)
@@ A.1 Case $\boldsymbol{\{P_{i\ell},D_{ijk},D_{jk\ell}\}}$
$[D_{ijk},[D_{jk\ell},P_{i\ell}]]$
$\qquad=[[P_{i\ell},D_{\ell jk}],D_{ijk}]=[P_{ik}P_{\ell j}-P_{\ell k}P_{ij},D_{ijk}]=[P_{\ell j}P_{ik}-P_{\ell k}P_{ij},D_{ijk}]$
$\qquad=[P_{\ell j},D_{jki}]P_{ik}-[P_{\ell k},D_{kij}]P_{ij}+P_{\ell j}[P_{ki},D_{kij}]-P_{\ell k}[P_{ij},D_{ijk}]$
$\qquad=(P_{i\ell}P_{jk}-P_{ij}P_{kl})P_{ik}-(P_{\ell j}P_{ik}-P_{jk}P_{i\ell})P_{ij}$
$\phantom{\qquad=}{}+P_{\ell j}(P_{ki}(P_{ij}-P_{jk})-2D_{ijk}-2P_{k}P_{ij}+2P_{i}P_{jk})$
$\phantom{\qquad=}{}-P_{\ell k}(P_{ij}(P_{jk}-P_{ki})-2D_{ijk}-2P_{i}P_{jk}+2P_{j}P_{ki})$
$\qquad=P_{i\ell}P_{jk}(P_{ik}+P_{ij})-(P_{ik}P_{j\ell}+P_{ij}P_{k\ell})P_{jk}$
$\phantom{\qquad=}{}+2(P_{k\ell}-P_{j\ell})D_{ijk}+2P_{i}(P_{j\ell}+P_{k\ell})P_{jk}-2P_{j}P_{k\ell}P_{ik}-2P_{k}P_{\ell j}P_{ij}.$ (A.2)